OpenStudy (maya):

integral from 1 to infinity of x^2*(e^-(2x))

OpenStudy (maya):

Explain with steps please

OpenStudy (anonymous):

1 to inf hell where'd you get that

OpenStudy (maya):

textbook

OpenStudy (maya):

I think you misunderstood the question. It is to the power of

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty} x^{2e^{2x}} \]right ?

OpenStudy (anonymous):

typo -2x not 2x .....

OpenStudy (maya):

\[\int\limits_{1}^{\infty} x ^{2} e ^{-2x}dx\]

OpenStudy (maya):

This is the correct one

OpenStudy (maya):

Power is -2x

OpenStudy (maya):

I dont know why the (-) sign does not show up

OpenStudy (anonymous):

or is it \[\int\limits_{1}^{\infty}x^2 \times e^{-2x}\] oh okay i got it ...its okay its a minus only ...i got it wring so many times lol

OpenStudy (anonymous):

you know by parts did you try it with that

OpenStudy (maya):

yes i tried that. I dont get the answer. I said let u= x^2, du=2xdx, dv=e^-2x and v=-1/2e^-2x

OpenStudy (anonymous):

if i try by parts then it will have me but how would you integrate it afterwards ...

OpenStudy (anonymous):

ignore the first sentence

OpenStudy (anonymous):

wait lemme gets some help ..

OpenStudy (anonymous):

see maya when i tried by parts then i get x^2 ..as one term ...but if i put inft it leads to undefined

OpenStudy (anonymous):

check your question again

OpenStudy (maya):

I get that too. But 1/inf = 0

OpenStudy (anonymous):

see after repeated by parts i get \[-\frac{x^2*e^{-2x}}{2} - \frac{x*e^{-2x}}{2} + \frac{e^{-2x}}{4}\] Now none of the expression gives a defined limit ...I could be wrong ..and I wish I'm wrong otherwise your question must have some errors

OpenStudy (maya):

Textbook answer is 5/(4*e^2)

OpenStudy (anonymous):

what ? i must be wrong the : (

OpenStudy (anonymous):

I'm slowly working it by parts but I'm a wee bit rusty on these.

OpenStudy (anonymous):

maya what if it is 2x*e^(x^2)

OpenStudy (maya):

nono Im sure that is the question.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+x^2*%28e^-%282x%29%29 I got the answers right

OpenStudy (anonymous):

trust me there is some problem with question

OpenStudy (anonymous):

I got the same answer as Maya. Give me a few minutes to type this beastie out.....

OpenStudy (anonymous):

These things were a beast in my Calc 2 class. They still are now :)

OpenStudy (anonymous):

hey there is no problem with your question http://www.wolframalpha.com/input/?i=+integrate+x^2*%28e^-%282x%29%29+from+1+to+infinity+ sorry for blaming your question

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}(x^{2}e^{-2x})dx\] Intergration by parts. \[Let u=x^{2}\] \[Let dv = e^{-2x}\] \[du=2xdx\] \[v=(-1/2)e^{-2x}\] Now we have: \[[(-1/2)x^{2}e^{-2x}]^{\infty}_{1} -\int\limits_{1}^{\infty}(2x(-1/2)e^{-2x})dx\] Solving the first part we get: \[(0-(1/2)e^{-2})\]

OpenStudy (anonymous):

-1 e^-2x(2x^2 + 2x + 1) ---x --------- 4 maybe infinity is not considered .....maybe if i put 1 then i get e^-2(5) --- 4

OpenStudy (anonymous):

which is your answer

OpenStudy (anonymous):

you have to ask someone else about it

OpenStudy (anonymous):

maya i got it here i'm posting the solution

OpenStudy (anonymous):

We now need to work the second part as another integration of parts: \[Let u = 2x\] \[du = 2dx\] \[Let dv = (-1/2)e^{-2x}\] \[v=(1/4)e^{-2x}\] Putting into the integration we get: \[-[[(1/2)xe^{-2x}]^{\infty}_{1}-\int\limits_{1}^{\infty}((1/2)e^{-2x})dx]\] Solving the left we get: \[-[(0-(1/2)e^{-2x})-\int\limits_{1}^{\infty}((1/2)e^{-2x})dx]\] Integrating the right we get: \[-[(1/2)e^{-2x} - [(1/4)e^{-2x}]^{\infty}_{1}\] Evaluating the right we get: \[-[(1/2)e^{-2x} - [0-(1/4)e^{-2}]]\] Putting it all back together you get: \[(1/2)e^{-2}+(1/2)e^{-2}+(1/4)e^{-2}=(5/4)e^{-2}=(5/(4e^{2}))\]

OpenStudy (anonymous):

now when we have to put infinity then we know integral is -1/4*e^-2x*(2x^2 + 2x +1) \[\lim_{x \rightarrow \infty } - \frac{e^{-2x}(2x^2 + 2x +1)}{4}\]

OpenStudy (anonymous):

which gives zero

OpenStudy (anonymous):

now all you've got is the second limit which is one so you get 5/4e^2

OpenStudy (anonymous):

BTW I made a typo in my first post. It should read: \[(0-(-1/2)e^{-2})\]

OpenStudy (anonymous):

That was a beast. I sure don't miss Calc 2. Nothing like page full solutions where one minus sign in the wrong spot is a wrong answer.

OpenStudy (anonymous):

Lol Osirisis

OpenStudy (maya):

Thank you :) Sorry I just got back to this problem again.

OpenStudy (anonymous):

Source: http://www.acalculator.com/math-calculators.html